OpenStudy (moongazer):

The point (3,2) is on a curve and at any point (x,y) on the curve, the tangent line has a slope equal to 2x-3. Find the equation of the curve.

OpenStudy (moongazer):

I only got to the part where the anti derivative is x^2 -3x +c

OpenStudy (amistre64):

you know the point 3,2 is on the line, it is called an initial condition.

OpenStudy (amistre64):

y = x^2 -3x + c 2 = 3^2 -3(3) + c, solve for c

OpenStudy (moongazer):

Thanks! :)

OpenStudy (amistre64):

yep